OpenStudy (anonymous):

Find the product of 543.1187 and 100 A. 5.431187 B. 54,311.87 C. 5,431.187 D. 54.31187 Can someone also explain how they came up with the answer so that I can understand? (do you move the decimal? If so, to the left or right? How many times you move it?

OpenStudy (anonymous):

Im stuck between two answers

OpenStudy (toman):

So you see that the decimal has 4 places to the right of the point, right?

OpenStudy (freckles):

here is examples \[9 \cdot \frac{1}{10000}=0.0009 \\ 9 \cdot \frac{1}{1000}=0.009 \\ 9 \cdot \frac{1}{100} =.09 \\ 9 \cdot \frac{1}{10}=.9 \\ 9 \cdot \frac{1}{1} =9 \\ 9 \cdot 10=90 \\ 9 \cdot 100=900 \\ 9 \cdot 1000=9000 \\ 9 \cdot 10000=90000\]

OpenStudy (freckles):

notice we are just moving the decimal point when we multiply by 10^(to some integer power)

OpenStudy (anonymous):

yes the decimal is over 4 that's why I thought it would need to be moved before you multiply

OpenStudy (anonymous):

Tge decimal doesn't move before you multiply. It moves whenever you multiply by a multiple of 10 which in this case is 100.

OpenStudy (anonymous):

The* sorry. Which answers are the ones you are stuck on?

OpenStudy (anonymous):

Im stuck between B and C

OpenStudy (anonymous):

I think it is B. It is just the ( , ) throwing me off. 54,311.87 or 5,431.187

OpenStudy (unklerhaukus):

100 has two zeros, so when multiplying by 100: move the decimal place two places to the right

OpenStudy (anonymous):

when I did that (moved decimal 2 places to the right) I got the 5431187 (no decimal in answer)

OpenStudy (unklerhaukus):